The perimeter of a field is 500 feet. Its width (shorter side) is 32 feet less than half its length (longer side). What is its width?
Answer by Taino(8)

Data given:
Perimeter of a rectangular field = 500 ft.
Length of the field = L
Width (W) of the field = 1/2*L - 32

Perimeter equation for a rectangle: P = 2*L + 2*W

500 = 2L + 2*(1/2*L -32)
500 = 2L + L - 64
500 = 3L - 64
500 + 64 = 3L

564 = 3L -----> L = 564/3 -----> L = 188 Ft.

Therefore, W = 1/2*L - 32 = 1/2*(188) - 32 = 94 - 32 = 62 Ft.
W = 62 Ft.

The width of the field is 62 feet.